One day as you are waiting for your friends at the (flat) soccer field you casually kick a

ball down the field. Amazingly you have kicked it at precisely 40◦ above horizontal at a

speed of 8.0m/s!! This is later verified through video analysis.

a) (8 points) How long is the ball in the air?

b) (9 points) How far from you is the ball when it lands? (Assume the answer to I) is 1.5

s if needed)


1
Expert's answer
2021-10-05T10:06:00-0400

a) "t=2v_0\sin\alpha/g=2\cdot8\cdot\sin40°/9.8=1.05\ (s)"


b) "l=v^2_0\sin2\alpha/g=8^2\cdot\sin(2\cdot40°)/9.8=6.43\ (m)"